Skip to content
free
K
1
2
practice
Home
Math
English
Science
Social Studies
Contact
Language
English
Español
Sign In
/
AP Courses
/
AP Biology
/
Multiplication Rule of Probability
Grade Level
AP Biology
APB.265 Multiplication Rule of Probability
0%
Skill Level
🔴 Hard
00:00
Progress to Next Level
25% to Medium
Easy
Medium
Hard
Mastery
100%
A student says that if two independent inheritance events must both occur, their probabilities should be added. Which calculation correctly revises the student's method for events with probabilities 1/4 and 3/4?
A
1/4 + 3/4 = 1
B
1/4 x 3/4 = 3/16
C
3/4 - 1/4 = 1/2
D
(1/4 + 3/4) / 2 = 1/2
Submit Answer